# One Look is Enough: A Novel Seamless Patchwise Refinement for Zero-Shot Monocular Depth Estimation Models on High-Resolution Images

### v1

@@ 5 Conclusion
Appendix A Architecture of the Fusion Module
The encoder takes $\mathbf{P}^{i}$, $\mathbf{D}_{\mathrm{f}}^{i}$, and $\mathsf{ROI}(\mathbf{D}_{\mathrm{c}})$ as inputs and produces a set of five intermediate features, denoted as $\mathbf{F}_{\mathrm{enc}}=\{\mathbf{f}_{\mathrm{enc},j}^{i}\}_{j=1}^{5}$ following [23]. Then, the fused feature map $\mathbf{F}_{\mathrm{fuse}}$ is obtained through the frequency fusion module ($\mathsf{FFM}$), defined as $\mathbf{F}_{\mathrm{fuse}}=\mathsf{FFM}(\mathsf{concat}(\mathbf{F}_{\mathrm{f}},\mathsf{ROI}(\mathbf{F}_{\mathrm{c}}))).$
@@ Appendix A Architecture of the Fusion Module
Subsequently, $\mathsf{concat}(\mathbf{F}_{\mathrm{fuse}},\mathsf{ROI}(\mathbf{F}_{\mathrm{c}}),\mathbf{F}_{\mathrm{enc}})$ is processed through two consecutive layers, each consisting of a $3\times 3$ convolution, batch normalization, and ReLU activation. Finally, the resulting feature is fed into the DPT decoder [31] to obtain the residual map $\mathbf{R}$.
@@
Architecture of the Frequency Fusion Module (FFM) To obtain accurate depth values from the coarse depth and preserve fine details from the fine depth, we design a Frequency Fusion Module (FFM) that effectively extracts and integrates edge information. We utilize Discrete Wavelet Transform (DWT) to decompose the input features into four frequency components: LL, LH, HL, and HH, which represent the low-frequency and high-frequency information. Each component is processed with its own dedicated convolution to capture scale-specific features. Finally, the components are recombined using the Inverse Discrete Wavelet Transform (IDWT), resulting in fused features that retain both global depth consistency and enhanced edge details. Overall process is described in Fig. 6-(b). To describe this process in more detail, we first decompose $\mathsf{ROI}(\mathbf{F}_{\mathrm{c}})$ and $\mathbf{F}_{\mathrm{f}}$ into four frequency sub-bands $\mathbf{X}$ ($\forall\mathbf{X}\in\{\mathbf{LL},\mathbf{LH},\mathbf{HL},\mathbf{HH}\}$) using $\mathsf{DWT}$. Each sub-band is then fused using a corresponding convolution $\text{Conv}_{\mathbf{X}}$. Finally, the fused feature map $\mathbf{F}_{\mathrm{fuse}}$ is obtained through $\mathsf{IDWT}$.
$\mathbf{X}_{\mathrm{c}},\mathbf{X}_{\mathrm{f}}=\mathsf{DWT}(\mathsf{ROI}(\mathbf{F}_{\mathrm{c}})),\mathsf{DWT}(\mathbf{F}_{\mathrm{f}})$ (8)
$\mathbf{X}_{\text{fuse}}=\text{Conv}_{\mathbf{X}}(\mathsf{concat}(\mathbf{X}_{\mathrm{c}},\mathbf{X}_{\mathrm{f}}))$ (9)
$\mathbf{F}_{\text{fuse}}=\mathsf{IDWT}(\mathbf{LL}_{\text{fuse}},\mathbf{LH}_{\text{fuse}},\mathbf{HL}_{\text{fuse}},\mathbf{HH}_{\text{fuse}})$ (10)
A.1 Ablation study of FFM
@@
To validate the effectiveness of the Frequency Fusion Module (FFM), we conduct an ablation study by replacing the FFM with a simple convolutional block consisting of Conv-ReLU-Conv layers. To ensure that any performance gain is not simply due to an increase in the number of parameters or FLOPs, we design the simple convolutional block to have more parameters and FLOPs than the FFM. This allows us to attribute the performance improvement to the design of FFM itself, rather than computational complexity. As shown in the Table 5, our method not only achieves the best performance in standard depth metrics, but also yields significant improvements in edge accuracy. Specifically, it achieves a 22.5% improvement on the DIS-5K dataset and a 16.9% improvement on the UHRSD dataset in the Boundary Recall (BR) metric. In addition, we observe a 4.6% improvement in the edge quality metric (D3R). It demonstrates that the proposed FFM effectively integrates edge information through the use of Discrete Wavelet Transform (DWT), which enables selective enhancement of high-frequency details without sacrificing global structure. This highlights the benefit of frequency-domain processing in depth refinement tasks.
Appendix B Qualitative Results
@@
Ablation Study In the ablation study (Sec. 4.4), we analyze the effect of Grouped Patch Consistency Training (GPCT) and Bias Free Masking (BFM) quantitatively. In this section, we analyze the effect of GPCT and BFM with qualitative results. As shown in the Fig. 7, the model trained without GPCT shows remarkable depth discontinuity problem on the grids. On the other hand, our PRO model trained with GPCT alleviates the depth discontinuity problem. Likewise, as shown in the Fig. 8 , the model trained without BFM exhibits artifacts on transparent surfaces, such as glass windows, as well as reflective surfaces like TV screens. In contrast, our model trained with BFM effectively refines only the edge regions while preventing artifacts on transparent objects.
Additional Qualitative Results
@@
We provide additional qualitative comparisons of BoostingDepth [26], PatchFusion [23], PatchRefiner [24], and PRO (Ours) on the UHRSD [45] dataset and on internet images (e.g., from Unsplash [^1] https://unsplash.com and Pexels [^2] https://www.pexels.com), as shown in Fig.9 and Fig.10.
@@
Figure 7: Qulitative comparisons of GPCT’s impact on ETH3D [35], Booster [29], and DIS-5k [28]. We compare PRO (Ours (d)) with the model trained without Grouped Patch Consistency Training (GPCT) (b). (b) and (d) represent the model index in Table 3. We also visualize the residuals to highlight the presence of artifacts more effectively. Zoom in for details.
@@
Figure 8: Qulitative comparisons of BFM’s impact on ETH3D [35], Booster [29], and DIS-5k [28]. We compare PRO (Ours (d)) with the model trained without Bias Free Masking (BFM) (c). (c) and (d) represent the model index in Table 3. We also visualize the residuals to highlight the presence of artifacts more effectively. Zoom in for details.
@@
Figure 9: Qulitative comparisons for patch-wise DE methods on UHRSD [45] and images from the internet. We compare PRO (Ours) with BoostingDepth [26], PatchFusion (PF) [23], and PatchRefiner (PR) [24]. The time displayed in the leftmost depth column represents the inference time. Black rectangle area represents the transparent object region. Black arrows indicate patch boundaries. Zoom in for details.
@@
Figure 10: Qulitative comparisons for patch-wise DE methods on images from the internet. We compare PRO (Ours) with BoostingDepth [26], PatchFusion (PF) [23], and PatchRefiner (PR) [24]. The time displayed in the top depth row represents the inference time. Black rectangle area represents the transparent object region. Black arrows indicate patch boundaries. Zoom in for details.

### v2

@@ Appendix A Extra Experiments
Appendix B Architecture of the Fusion Module
@@
The encoder takes $\mathbf{P}^{i}$, $\mathbf{D}_{\mathrm{f}}^{i}$, and $\mathsf{ROI}(\mathbf{D}_{\mathrm{c}},\mathbf{P}^{i})$ as inputs and produces a set of five intermediate features, denoted as $\mathbf{F}_{\mathrm{enc}}=\{\mathbf{f}_{\mathrm{enc},j}^{i}\}_{j=1}^{5}$ following [23]. Then, the fused feature map $\mathbf{F}_{\mathrm{fuse}}^{i}$ is obtained through the frequency fusion module ($\mathsf{FFM}$), defined as $\mathbf{F}_{\mathrm{fuse}}^{i}=\mathsf{FFM}(\mathsf{concat}(\mathbf{F}_{\mathrm{f}}^{i},\mathsf{ROI}(\mathbf{F}_{\mathrm{c}},\mathbf{P}^{i}))).$ Subsequently, $\mathsf{concat}(\mathbf{F}_{\mathrm{fuse}}^{i},\mathsf{ROI}(\mathbf{F}_{\mathrm{c}},\mathbf{P}^{i}),\mathbf{F}_{\mathrm{enc}})$ is processed through two consecutive layers, each consisting of a $3\times 3$ convolution, batch normalization, and ReLU activation. Finally, the resulting feature is fed into the DPT decoder [31] to obtain the residual map $\mathbf{R}$.
@@
Architecture of the Frequency Fusion Module (FFM) To obtain accurate depth values from the coarse depth and preserve fine details from the fine depth, we design a Frequency Fusion Module (FFM) that effectively extracts and integrates edge information. We utilize Discrete Wavelet Transform (DWT) to decompose the input features into four frequency components: LL, LH, HL, and HH, which represent the low-frequency and high-frequency information. Each component is processed with its own dedicated convolution to capture scale-specific features. Finally, the components are recombined using the Inverse Discrete Wavelet Transform (IDWT), resulting in fused features that retain both global depth consistency and enhanced edge details. Overall process is described in Fig. 8-(b). To describe this process in more detail, we first decompose $\mathsf{ROI}(\mathbf{F}_{\mathrm{c}},\mathbf{P}^{i})$ and $\mathbf{F}_{\mathrm{f}}^{i}$ into four frequency sub-bands $\mathbf{X}$ ($\forall\mathbf{X}\in\{\mathbf{LL},\mathbf{LH},\mathbf{HL},\mathbf{HH}\}$) using $\mathsf{DWT}$. Each sub-band is then fused using a corresponding convolution $\text{Conv}_{\mathbf{X}}$. Finally, the fused feature map $\mathbf{F}_{\mathrm{fuse}}^{i}$ is obtained through $\mathsf{IDWT}$.
$\mathbf{X}_{\mathrm{c}}^{i},\mathbf{X}_{\mathrm{f}}^{i}=\mathsf{DWT}(\mathsf{ROI}(\mathbf{F}_{\mathrm{c}},\mathbf{P}^{i})),\mathsf{DWT}(\mathbf{F}_{\mathrm{f}}^{i})$ (8)
$\mathbf{X}_{\text{fuse}}^{i}=\text{Conv}_{\mathbf{X}}(\mathsf{concat}(\mathbf{X}_{\mathrm{c}}^{i},\mathbf{X}_{\mathrm{f}}^{i}))$ (9)
$\mathbf{F}_{\text{fuse}}^{i}=\mathsf{IDWT}(\mathbf{LL}_{\text{fuse}}^{i},\mathbf{LH}_{\text{fuse}}^{i},\mathbf{HL}_{\text{fuse}}^{i},\mathbf{HH}_{\text{fuse}}^{i})$ (10)
B.1 Ablation study of FFM
@@
To validate the effectiveness of the Frequency Fusion Module (FFM), we conduct an ablation study by replacing the FFM with a simple convolutional block consisting of Conv-ReLU-Conv layers. To ensure that any performance gain is not simply due to an increase in the number of parameters or FLOPs, we design the simple convolutional block to have more parameters and FLOPs than the FFM. This allows us to attribute the performance improvement to the design of FFM itself, rather than computational complexity. As shown in the Table 9, our method not only achieves the best performance in standard depth metrics, but also yields significant improvements in edge accuracy. Specifically, it achieves a 22.5% improvement on the DIS-5K dataset and a 16.9% improvement on the UHRSD dataset in the Boundary Recall (BR) metric. In addition, we observe a 4.6% improvement in the edge quality metric ($\mathrm{D}^{3}\mathrm{R}$). It demonstrates that the proposed FFM effectively integrates edge information through the use of Discrete Wavelet Transform (DWT), which enables selective enhancement of high-frequency details without sacrificing global structure. This highlights the benefit of frequency-domain processing in depth refinement tasks.
Appendix C Additional Implementation Details
Details in training. Before training, we first filter out unnecessary training samples based on the unreliable mask, $\textbf{M}_{\mathrm{unreliable}}$, described in Section 3.3. Specifically, we load each $\textbf{M}_{\mathrm{unreliable}}$ using NumPy and discard the corresponding training sample if the mean value of $\textbf{M}_{\mathrm{unreliable}}$ exceeds 0.5, indicating excessive unreliable regions. During training, we randomly crop a region (e.g., 846$\times$1505) from a 2160$\times$3840 input to generate 2$\times$2 overlapping patches (540$\times$960 each), which are resized to 518$\times$518 for refinement. The overlap size (234$\times$415) corresponds to $\frac{224}{518}\approx 43\%$ of the patch dimensions, where 224 is the empirically chosen overlap value from Section 4.4.
Details in inference. At inference, the input is divided into a 4$\times$4 non-overlapping grid, and the refined 518$\times$518 patches are reassembled into a 2072$\times$2072 depth map. The final depth map is then upsampled to the original resolution via bilinear interpolation. This inference procedure is consistent with PFP=16 and PRP=16. As demonstrated in Table 7, the number of patches can be flexibly adjusted depending on the inference setting.
Appendix D Qualitative Results
@@
Ablation Study In the ablation study (Section 4.4), we analyze the effect of Grouped Patch Consistency Training (GPCT) and Bias Free Masking (BFM) quantitatively. In this section, we analyze the effect of GPCT and BFM with qualitative results. As shown in the Fig. 9, the model trained without GPCT shows remarkable depth discontinuity problem on the grids. On the other hand, our PRO model trained with GPCT alleviates the depth discontinuity problem. Likewise, as shown in the Fig. 10 , the model trained without BFM exhibits artifacts on transparent surfaces, such as glass windows, as well as reflective surfaces like TV screens. In contrast, our model trained with BFM effectively refines only the edge regions while preventing artifacts on transparent objects.
@@
Additional Qualitative Results We provide additional qualitative comparisons of BoostingDepth [26], PatchFusion [23], PatchRefiner [24], and PRO (Ours) on the UHRSD [45] dataset and on internet images (e.g., from Unsplash [^1] https://unsplash.com and Pexels [^2] https://www.pexels.com), as shown in Fig.11 and Fig.12.
@@
Figure 9: Qualitative comparisons of GPCT’s impact on ETH3D [35], Booster [29], and DIS-5k [28]. We compare PRO (Ours (d)) with the model trained without Grouped Patch Consistency Training (GPCT) (b). (b) and (d) represent the model index in Table 3. We also visualize the residuals to highlight the presence of artifacts more effectively. Zoom in for details.
@@
Figure 10: Qualitative comparisons of BFM’s impact on ETH3D [35], Booster [29], and DIS-5k [28]. We compare PRO (Ours (d)) with the model trained without Bias Free Masking (BFM) (c). (c) and (d) represent the model index in Table 3. We also visualize the residuals to highlight the presence of artifacts more effectively. Zoom in for details.
@@
Figure 11: Qualitative comparisons for patch-wise DE methods on UHRSD [45] and images from the internet. We compare PRO (Ours) with BoostingDepth [26], PatchFusion (PF) [23], and PatchRefiner (PR) [24]. The time displayed in the leftmost depth column represents the inference time. Black rectangle area represents the transparent object region. Black arrows indicate patch boundaries. Zoom in for details.
@@
Figure 12: Qualitative comparisons for patch-wise DE methods on images from the internet. We compare PRO (Ours) with BoostingDepth [26], PatchFusion (PF) [23], and PatchRefiner (PR) [24]. The time displayed in the top depth row represents the inference time. Black rectangle area represents the transparent object region. Black arrows indicate patch boundaries. Zoom in for details.